solve this system of linear equations. separate the x- and the y-values with a comma?
6x=-58-11y
-4x=2-11y

Answer by ankor@dixie-net.com(22740)

6x=-58-11y
-4x=2-11y
Rearrange to use elimination
6x + 11y = -58
-4x+ 11y = 2
-----------------Subtraction eliminates y, find x
10x + 0- = -60
x = -60/10
x = -6
Find y using the first equation
6(-6) = -58 - 11y
11y = -58 + 36
11y = -22
y = -22/11
y = -2
:
:
Check solution in the 2nd equation
-4(-6) = 2 - 11(-2)
+24 = 2 + 22
:
x,y: -6.-2